Im sorry but can you help me with the other question

Angelo thinks that x^2=(x-y)^2 for all values of x and y. which pair of values disproves his theory?
A x=2 and y=0
B x=1 and y=1
C x=0 and y=1
D x=0 and y=0

Answer by edjones(8007)

B) 1^2=(1-1)^2
1=0
.
C) 0^2=(0-1)^2
0=1
.
Both are wrong.
.
Ed